OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2.

OpenStudy (anonymous):

do you know the general slope intercept form of a line?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

@mathmaster, i told u the answer

OpenStudy (anonymous):

my computer spazed out lol

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq.htm

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

y=1/2 x+c here, y=2, x=4 so, 2=1/2 *4+c so, c=0 so, the slope-intercept form is : y=1/2 x

OpenStudy (anonymous):

well, the general form of a line represented in the slope intercept form is y = mx+b where m is the slope and b is the intercept. you know m = 1/2 and that the line passes through the point (4,2), which means that x = 4 y = 2 will solve the equation. find b.

OpenStudy (anonymous):

thnks guys